A sample of rubbing alcohol contains 142 g of isopropyl alcohol (C2H7OH) and 58 g of water. What are the mole fractions of alcohol and water?

1
Expert's answer
2021-02-23T04:30:22-0500

moles of isopropyl alcohol = 142/60.01 = 2.36moles

moles of water = 58/18 = 3.22moles

For isopropyl alcohol C3H7OH


"\frac{2.36 moles C_3H_7OH}{2.36 moles C_3H_7OH+3.22 moles H_2O}" = 0.423


For water


"\frac{3.22 moles H_2O}{2.36 moles C_3H_7OH+3.22 moles of H_2O}" = 0.577


Mole fraction of C3H7OH is 0.423 and H2O is 0.577